### 20 Mar 2026 v1

# The dusty envelopes of asymptotic giant branch stars with ultraviolet excesses

Abstract
Context. Roughly spherical envelopes around asymptotic giant branch (AGB) stars transform into the highly-asymmetric morphologies observed in planetary nebulae. The complex processes suffered in this metamorphosis are not yet completely understood. However, binarity emerges as a strong shaping factor, although the identification of binary companions in AGB stars is observationally challenging. The presence of ultraviolet (UV) excesses in AGB stars has been suggested as a potential indicator of binarity.
Aims. In a first study, we characterised the properties of the gas component in the circumstellar envelopes surrounding a sample of 29 AGB stars with UV excesses. Now we intend to complement this information with an analysis of the dust component and compare the estimated parameters with those previously inferred from larger samples of AGB stars.
Methods. We modelled the spectral energy distributions of the sample using dust radiative transfer models. In some cases, we complemented the analysis with Herschel/PACS radial surface brightness profiles.
Results. We derived mass-loss rates and gas-to-dust ratios, which are in the typical ranges for AGB stars. We found that the stellar and mass-loss parameters follow similar trends than those presented in the literature. There is an anticorrelation between the gas-to-dust ratio and the UV emission, although it is weaker than its correlations with pulsation and mass-loss. We also estimated the dust attenuation produced by the dust at UV wavelengths and describe its effects on the intrinsic UV emission.
Conclusions. Stellar and mass-loss parameters of UV emitting AGB stars follow similar trends as found for larger samples of AGB stars. High-angular resolution observations are required to explore the dust forming regions and identify the presence of stellar companions. Circumstellar dust attenuation might play a dominant role in the observed UV emission, and needs to be accounted to estimate the intrinsic UV emission.
Key Words.:
Stars: AGB and post-AGB – circumstellar matter – mass-loss – Ultraviolet: stars – dust, extinction1 Introduction
When stars with low and intermediate masses (i.e. with masses 1-8 ) leave the main-sequence, they evolve through the Hertzsprung–Russell diagram (HRD) increasing their luminosities while lowering their effective temperatures. They eventually reach the Asymptotic Giant Branch (AGB). At this stage their huge luminosities (from hundreds to thousands of solar luminosities) and intense stellar pulsation induce high mass-loss rates (- yr-1). The large amount of material transferred by these stars to the interstellar medium (ISM) causes AGB sources to become a major source of galactic chemical enrichment (see Tielens, 2005; Ferrarotti & Gail, 2006).
During this process, the stellar material is expelled at low velocities (5-30 km s-1) and accumulates surrounding the stars, where the low temperatures allow the formation of molecular gas and dust in the vicinity of the star (see Höfner & Olofsson, 2018). This process leads to the formation of dense Circumstellar Envelopes (CSEs) that display (roughly) spherical shapes Castro-Carrizo et al. (2010). In addition, AGB CSEs show a very rich molecular and dust chemistry (for a recent overview see Agúndez et al., 2020), which is mainly determined by the carbon-to-oxygen ratio (C/O). The chemistry in O-rich AGB stars (C/O1) is dominated by O-bearing molecules (e.g. , SiO) and related dust (silicates, oxides), whereas the chemistry in C-rich AGB stars (C/O1) is dominated by C-bearing molecules (e.g. , HCN) and related dust (carbon, carbides).
When these stars leave the AGB a dramatic transformation is observed in the shape of their CSEs, which might metamorphose into the large variety of shapes commonly found in Planetary Nebulae (PNe) and pre-PNe. They include highly elliptical, bipolar, or multipolar structures (Sahai & Trauger, 1998; Ueta et al., 2000; Sahai et al., 2007, 2011a; Stanghellini et al., 2016) that greatly contrast the shapes of AGB CSEs (see Balick & Frank, 2002). Moreover, during these post-AGB stages high-velocity outflows develop and CSEs become optically thinner allowing a high degree of ionisation by the central star (see e.g. Kwitter & Henry, 2022).
Even though it is still not well understood which processes lead to this morphological and kinematic transformation, binarity is expected to play a major role in shaping the CSEs (see e.g. Nordhaus & Blackman, 2006; De Marco, 2009). Binarity should be common in AGB stars since it is found to be common on the main sequence (see Duquennoy & Mayor, 1991) as well as during the post-AGB phase(see Miszalski et al., 2009). However, the direct identification of stellar companions of AGB stars remains a major challenge due to their huge luminosity, the extinction produced by their CSEs, and the intrinsic variability (in both photometry and radial velocities) induced by stellar pulsations.
Recently, high-angular resolution observations have identified the formation of binary-induced structures (e.g. spirals and rotating discs) on small scales in some AGB stars (see e.g. Decin et al., 2020, and references therein). These structures, likely shaped by stellar companions, can serve as initial seeds for the development of the complex morphologies found in PNe.
On the other hand, Sahai et al. (2008) discovered a subclass of AGB stars that show an intense ultraviolet (UV) emission (hereafter uvAGBs), orders of magnitude higher than the expected. These UV excesses can be explained by the presence of hot stellar companions and accretion processes (see e.g. Sahai et al., 2008; Ortiz & Guerrero, 2016), although other phenomena like chromospheric activity cannot be ruled out (see Montez et al., 2017). Moreover, complementary observational efforts have found a large UV variability (Sahai et al., 2011b, 2016, 2018), continuum-dominated UV spectra (Ortiz et al., 2019) and the detection of X-ray emission (Ramstedt et al., 2012; Sahai et al., 2015; Ortiz & Guerrero, 2021) in some of these stars. In these cases, the most likely explanation is the presence of interacting stellar companions and accretion processes.
Recently, Sahai et al. (2022) performed a series of modelling studies and proposed that the origin of UV excesses can be related with the ratio between FUV and NUV fluxes (). UV excesses with 0.06 are likely produced by intense accretion, whereas 0.06 can be produced by chromospheric activity and/or less intense accretion processes.
The first characterisation of the CSEs of uvAGBs as a class was performed by Alonso-Hernández et al. (2024), who studied their mass-loss properties based on CO rotational emission lines. One of their results was a statistically lower CO intensity against the 60 m flux in comparison with large samples of AGB stars, indicating lower amounts of molecular gas with respect to dust (i.e. lower gas-to-dust ratios).
In this paper we present the first characterisation of the dust component of the CSEs, based on the modelling of Spectral Energy Distributions (SEDs), in uvAGBs and explore the effects of dust opacity on the observed UV emission. In Sect. 2 we describe our sample and observations. In Sect. 3 we describe the modelling procedure, the comparison with the observational data and the space of parameters explored. In Sect. 4 we describe the main results obtained in this study, which are later discussed in Sect. 5. Finally, we summarised our main conclusions in Sect. 6.
2 Observational data sets
In this study, we specifically focus on the sample of 29 UV emitting AGB stars presented in Alonso-Hernández et al. (2024), which is based on a CO emission line survey of AGB stars with UV emission detected by the Galaxy Evolution Explorer (GALEX, Martin et al., 2005). The equatorial coordinates, distances and complementary information about the sources can be found in Table 1 of Alonso-Hernández et al. (2024).
We based our analysis on three kinds of archival data, which are classified as follows: (i) photometric fluxes obtained from general catalogues and surveys, (ii) optical and infrared spectra, and (iii) Herschel/PACS images from which we extracted the photometric fluxes and radial surface brightness profiles.
2.1 Photometry
The SEDs were built using previously calibrated photometric fluxes that cover the optical to far infrared (FIR) range (0.5-200 m) with the most reliable fluxes using the VizieR Photometry viewer tool available at VizieR database (Ochsenbein et al., 2000). In the optical range, we selected photometry corresponding to the Johnson B and V, Gaia , G and filters. In the infrared range, we selected the 2MASS J, H and K, Infrared Astronomical Satellite (IRAS, Neugebauer et al., 1984) and AKARI (Doi et al., 2015) photometry (quality flag 2 or 3 for detections and 1 for upper limits respectively). The available WISE observations were not used due to saturation, as it is common in most galactic AGB stars (see e.g. Suh, 2018).
We noted the presence of near-field objects in the vicinity of some of our targets. However, in all the cases our AGB stars are significantly brighter than the surrounding objects and the effect of pollution was estimated, using the observations with highest angular resolution (e.g. Gaia in the optical and Herschel/PACS in the FIR, see also Sect. 2.3), to be around or less than 1% of the fluxes (i.e. lower than total flux uncertainties).
When multi-epoch photometric values were available, we used the average fluxes. The flux uncertainties were estimated as the sum of their standard deviation and the average of their individual uncertainties. These uncertainties are dominated by the standard deviation, as they typically show a significant scattering related to the flux variability induced by the AGB pulsations.
2.2 Spectra
We gathered some large wavelength coverage spectra in order to obtain a more complete view of these objects. In particular, we mainly used Gaia DR3 XP (see De Angeli et al., 2023) as optical spectra (covering 340-1020 nm with R20-200) for 25 sources and IRAS Low Resolution Spectra (LRS) from the extended LRS atlas111Publicly available on https://users.physics.unc.edu/~gcsloan/library/lrsatlas (see Sloan et al., 2025) as infrared spectra (covering 7.7-22.7 m with R20-60) for the whole sample.
Moreover, we used Infrared Space Observatory (ISO, Kessler et al., 1996, 2003) Short Wavelength Spectrometer (SWS; de Graauw et al., 1996; Leech et al., 2003) from the SWS atlas222Publicly available on https://users.physics.unc.edu/~gcsloan/library/swsatlas (see Sloan et al., 2003b) for 3 sources and Long Wavelength Spectrometer (LWS; Clegg et al., 1996; Gry et al., 2003) from the ISO archive for 1 source. We used the available ISO SWS (covering 2.4-45.2 m with R1000-2000) and ISO LWS (covering 43-197 m with R150-200) when available as infrared spectra instead of IRAS LRS due to their larger wavelength coverages and higher spectral resolutions.
2.3 Herschel/PACS imaging
We used archival observations performed with the Photodetector Array Camera and Spectrometer (PACS; Poglitsch et al., 2010) on board Herschel Space Observatory (Pilbratt et al., 2010), in particular photometric images in its three bands (“BLUE” at 70 m, “GREEN” at 100 m, “RED” at 160 m) for seventeen sources. Four sources have available level 2.5 “scan” data, we used the images combined with the UNIMAP method (Piazzo et al., 2015) for VY UMa and with the JSCANAM method (Graciá-Carpio et al., 2015) for V Eri, EY Hya and IN Hya. Moreover, thirteen sources have level 2.0 data, we combined the available “scan”, as a first option, or “chop-nod”, as a second option, images with standard Herschel Interactive Processing Environment (HIPE, Ott, 2010) scripts.
In first place, we checked whether the angular sizes of our targets are well-described by Herschel/PACS Point Spread Functions (PSFs). For this purpose, we estimated the PACS radial surface brightness profiles and compared them with those of the PSFs. We found this case in most of our targets (“point sources”), whereas in four sources (RU Her, SV Peg, T Dra, and V Eri) the angular size are slightly larger than the PSFs (“semi-extended sources”, see appendix A). Furthermore, VY UMa presents an extended detached shell with an extension of 35-70″ (previously reported by Cox et al., 2012; van Marle et al., 2014) that dominates the integrated FIR emission (see appendix B).
We performed open aperture photometry removing near-field sources in the Herschel/PACS image to estimate accurately their fluxes. For the “point sources”, the photometric aperture radii were the standard values for point sources (12″, 20″ and 25″ in the BLUE and GREEN bands and 22″, 24″ and 28″ in the RED band), whereas for “semi-extended sources” we applied photometric aperture radii of 22″, 24″ and 28″ respectively.
In the case of VY UMa, we used aperture sizes of 12″ for the BLUE and GREEN bands as well as 22″ for the RED band, subtracted the surface brightness at the centre of the detached shell, and estimated the sky background in clear regions. We noted a large discrepancy with respect to the fluxes of IRAS and AKARI at similar wavelengths because their larger beams include, at least partially, the detached shell. We additionally estimated the emission of the detached shell, with aperture sizes of 70″ in the three bands, and propose a model to fit the extended emission of this detached shell (see appendix B).
The relative flux uncertainties from the individual and combined Herschel/PACS images can be lower than . However, the flux variability between different epochs, even considering the scarcity of multi-epoch images and the sparse temporal coverage for most sources, is significantly larger (in some cases at least 5-20%). Therefore, the instrumental uncertainties underestimate the intrinsic variability of AGB stars on the far-infrared (this also can apply to IRAS and AKARI photometry). The employed Herschel/PACS observations and their respective estimated fluxes are summarised in appendix C.
3 Analysis: dust emission model
The circumstellar dust produces absorption, scattering, and thermal emission across the whole electromagnetic spectra, leading to changes in the overall shape of SEDs. Therefore, the comparison between the observed SEDs and those expected for the photospheres of AGB stars allows to characterise simultaneously the main physical parameters of both the stellar (e.g. the effective temperature, luminosity and roughly the C/O ratio) and dust (e.g. optical depth, dust density and temperature) components of the CSE. A broad wavelength coverage is particularly relevant to this analysis, although complementary infrared spectroscopy can be useful to constrain the dust composition by fitting the spectral features produced by the solid-state bands from the dust.
3.1 Description of the DUSTY envelope models
We modelled the optical to far-infrared SED of our 29 targets with the 1D-radiative transfer code DUSTY (Ivezic et al., 1999), which creates synthetic spectra based on the emission originated by a central source surrounded by spherical dust layers. The main six variables for DUSTY models are: (i) input stellar spectra, (ii) dust chemical composition, (iii) grain size distribution, (iv) temperature at the inner radius (), (v) density distribution and (vi) optical depth.
We note that uvAGBs are binary candidates in which it is expected the presence of accreting stellar companions, which should be located inside the dust formation zone, that can produce shaping mechanisms, mass transfer and an anisotropic radiation field. These phenomena may lead to significant deviations from spherical symmetry in their CSEs, with a more pronounced effect in the innermost regions. Nevertheless, in the absence of spatially-resolved observations, we consider spherical symmetry to be a good approximation of the bulk of the CSE. We discuss possible effects of this approximation in Sect. 5.
The stellar component (input spectra) used as an input for DUSTY was included with the COMARCS stellar atmosphere library333Publicly available on http://stev.oapd.inaf.it/atm/index.html (see Aringer et al. (2016) for O-rich stars and Aringer et al. (2019) for C-rich stars). These stellar models mainly depend on the stellar effective temperature () and the ratio between carbon and oxygen atomic abundances.
Given the lack of accurate measurements, we selected COMARCS models with standard abundances of C/O0.55 for oxygen-rich stars and C/O1.40 for carbon-rich stars and solar metallicity (e.g. Agúndez et al., 2020). We remark that a variation of the C/O ratio only affects significantly the stellar spectra, in terms of SED modelling, at values near C/O1.00 due to the drastic change from O-rich to C-rich chemistry (see Aringer et al., 2016, 2019).
The dust composition is a key parameter that mainly affects the shape of the SED in the infrared, producing different and characteristic broad spectral features and modifying the shape of the continuum emission. For the O-rich AGB stars, we assumed a mixture of silicates (from Draine & Lee, 1984), alumina (Al2O3, from Begemann et al., 1997) and iron oxide (FeO, from Henning et al., 1995). Whereas, for the C-rich AGB stars we assumed a mixture of amorphous carbon (from Hanner, 1988), silicon carbide (SiC, from Pegourie, 1988) and graphite (from Draine & Lee, 1984).
We performed a basic characterisation of the dust composition only for the three targets with available ISO SWS spectra. For this purpose, we performed an iterative process varying the relative dust abundances in 5% steps until we found solutions that fitted the overall SEDs (see Sect. 3.2) and the infrared spectral features. We obtain the following best-fit grain compositions: RW Boo (60% silicates, 30% O3 and 10% FeO), SV Peg (50% silicates, 30% O3 and 20% FeO) and T Dra (80% amorphous carbon, 15% SiC and 5% graphite), which reproduce the spectral features associated to the assumed dust compounds, although, we noted the presence of other small features.
We assumed similar compositions for the rest of targets based on their chemical type (for O-rich 50% silicates, 40% Al2O3 and 10% FeO and for C-rich 80% amorphous carbon, 5% graphite and 15% SiC). These compositions are in agreement with those expected according to the chemical type and mass-loss rates of our targets (see e.g. Waters, 2011). However, we acknowledge that dust composition is a major source of uncertainty due to the presence of spectral features with unknown carriers (see e.g. Sloan et al., 2003a), featureless dust components (e.g. metallic iron, see Kemper et al., 2002), and large degeneracies between the different spectral features and with the overall shape of the SED (see e.g. Speck et al., 2000; Jones et al., 2014).
On the other hand, the infrared region of the SEDs is also quite dependent of the grain size distribution (see e.g. Ysard et al., 2018), although it is very complicated to constrain empirically. In this study we used the standard Mathis-Rumpl-Nordsieck (MRN) distribution (Mathis et al., 1977), which is the most common distribution in circumstellar and interstellar dust studies (see references in Mathis et al., 1977), described as:
where q3.5, =0.005 m, and =0.25 m are respectively the exponent, the minimum and the maximum. These are the standard MNR values of the dust grain size distribution.
The temperature at the inner radius of the envelope indicates the maximum temperature of the dust affecting, therefore, its thermal emission. This parameter is completely correlated with the inner radius of the envelope (), which is also a DUSTY output, through the temperature distribution (Ivezic et al., 1999).
The density distribution is assumed to be a power-law:
where is the outer radius, and , the index of the power-law, is a free parameter. In the case of a spherical wind expanding at a constant velocity, the density distribution is described by an index =2 (see Ivezic & Elitzur, 1995). The ratio = indicates the extension of the envelope.
Finally, the radial optical depth, , at a fiducial wavelength of 550 nm, is used to parametrise the optical depth of the CSE across the entire wavelength range covered in the model. We used the optical depths at the central wavelengths of the two GALEX bands in order to estimate the effect of dust attenuation in the observed UV emission (see appendix D).
3.2 Parameter space and SED fitting
We performed a systematic search for the best-fit models with five free parameters: , , , , and . To determine the best-fit parameters of the CSEs of the fit we perform a reduced minimisation, with described as:
where is the photometric flux of each observation, is its associated flux uncertainty, is the synthetic photometry estimated from the model, N is the number of observations and is the number of free parameters (in our case 5), and is the number of degrees of freedom of the fitting. Spectral observations were not used to estimate the because they contain gas spectral lines that were not included in our SED modelling.
We estimated the synthetic photometry in each filter convolving the synthetic spectra produced by DUSTY with their respective transmission curves, which are provided by the SVO Filter Profile Service “Carlos Rodrigo” (Rodrigo et al., 2012; Rodrigo & Solano, 2020; Rodrigo et al., 2024). In addition, we applied ISM extinction correction to the SEDs using the extinction law presented in Gordon et al. (2023) and the E(B-V) values from Alonso-Hernández et al. (2024). Finally, we scaled all the synthetic photometry by the coefficient that minimised the .
Particular fluxes with very low relative uncertainties might induce model overfitting in certain spectral regions, leading to wrong best-fit solutions. As mentioned in Sect. 2, far infrared single-epoch observations usually have very low flux uncertainties. However, a large dispersion of the photometric points due to the stellar pulsations is expected (see e.g. Onaka et al., 1999; Groenewegen, 2022; Tachibana et al., 2023), and also at larger wavelengths (see e.g. Jenness et al., 2002; Dehaes et al., 2007).
Our simple study on the multi-epoch Herschel/PACS observations (see Sect. 2.3), indicates that variability-induced uncertainties should be at least around 5%. Therefore, we impose a minimum of 5% uncertainties in all photometric fluxes:
where is the uncertainty estimated as the sum of the instrumental error and the standard deviation when multi-epoch measurements are available. This conservative limit prevents from model over-fitting and provides reasonable best-fit solutions.
Moreover, we noted that a few sources (R LMi, R UMa and Z Cnc) show a large difference between the fluxes obtained from Herschel/PACS and those from IRAS and AKARI (see Fig. 1). We noted that, in these cases, IRAS and AKARI fluxes are systematically higher, indicating pollution from an irregular sky brightness due to their more limited angular resolution. This large difference in a short wavelength range produces a large increase of and complicates the overall reliability of the fitting. Therefore, we did not include these fluxes in the analysis.
To explore the parameter space, we employed a simple fitting procedure, searching for an initial guest solution by visual inspection. Once we arrived to a good solution, we performed a numerical fitting with a grid of the five free parameters and chose the solution with the lowest as the best-fit solution.
We used linear spaces for , , n and , and a logarithmic space for Y. The accuracy and ranges of the were limited by the availability of the stellar photospheric models, whereas for the rest of parameters they were determined considering the steepness of the distribution in each case. A discussion about the correlations between the different parameters and their associated uncertainties is provided in Sect. 4.1.
4 Results
We present the main results from our characterisation of the dust component of the CSEs of our sample. We describe the best-fit models obtained from the SED modelling, derive additional parameters, and explore correlations between the parameters.
4.1 Best-fit solutions and uncertainties
We find that the SEDs of most of our targets (21/29) show an infrared excess and require a dusty CSE (see Figs. 1 and table 1); we derive additional parameters for these sources in Sects. 4.2 and 4.3. The SEDs of some sources (8/29) can be fitted using only stellar photospheric models (see Fig. 2 and table 2 implying a negligible dust component, presumably because they have not yet experienced substantial mass-loss. We refer to such sources as “naked” (as defined by Sloan & Price, 1995). Furthermore, we note that all the sources detected in CO by Alonso-Hernández et al. (2024) have infrared excess, whereas not all dusty AGB stars were detected in CO.
We highlight that all the models fit the SEDs with 1.0 (ranging 0.6-7.2) and show good agreements with the available spectra, for both dusty and naked AGB stars. Our best-fit DUSTY models reproduce satisfactorily the radial surface brightness profiles of semi-extended sources (see Appendix A), including the detached shell of VY UMa (see appendix B).
Fig.3 shows the distributions of the best-fit values for , , and . We find that dusty uvAGBs have in the range 2600-3500 K, with similar values for both O-rich and C-rich AGB stars, although we acknowledge that the sample of C-rich is very limited. The “naked” stars are located at the highest temperatures of the distribution (3400-3500 K) and also present low luminosities (800-3000 ), indicating that they might be in the early-AGB phase or still on the Red Giant Branch (see also Alonso-Hernández et al., 2024).
All the sources have values of lower than dust sublimation temperatures (typically 1400-1500 K, see Agúndez et al., 2020). In some cases, is rather low (600-1000 K, see Fig.3 middle-left), which indicates a relative lack of warm dust in the immediate vicinity of the star.
The distribution of is around the expected values of 2 for radiatively driven winds (see e.g. Ivezic & Elitzur, 1997), although in some cases they are lower (see Fig.3 middle-right). These low values might indicate deviations from smooth outflows (we resume this point in Sect. 4.4).
Most of the sources presented in this study have optically-thin envelopes at the fiducial wavelength 1, although a few are optically-thick (1, see Fig.3 right). This clustering at low can be expected because uvAGBs with low optical depths have a higher chance of being detected via their UV emission due to a low attenuation (see Sect. 5 and Appendix D).
We found that, in some cases, is rather uncertain because the wavelength coverage of the observations does not extend to the sufficiently long wavelength that is required to probe the coldest dust located in the outermost regions of the envelopes (see also Heras & Hony, 2005).
We performed a sensitivity analysis to determine the uncertainties and correlations of the free parameters. The uncertainties were estimated from the log-likelihood confidence intervals and show asymmetric distributions. The uncertainties of the parameters are also affected by the spectral variability of AGB stars. In the case of targets without a dust component, is the only parameter correctly estimated as converges to zero in all cases (0.01) and the other parameters do not influence the fit.
We found that the free parameters are degenerate and their uncertainties systematically underestimate second order effects, that is the parameters compensate the effects on the spectral shape between them. Some of these correlations display a characteristic “banana” shape indicating non-linearity (we present a detailed discussion about these correlations in appendix E).
[b] Source log() L Mdust (K) (K) () (cm) (cm) ( ) ( ) EY Hya 3000 1350 1.86 0.35 3.8 5200800 7.9 50 29 19 310 FH Vir 3400 1500 1.75 0.05 4.4 2100300 3.3 83 13 – – IN Hya 3500 1450 1.78 0.04 4.0 1700200 3.3 33 2.4 – – R LMi 3100 1250 2.03 1.6 4.6 3300500 4.8 190 48 5.2 810 R UMa 3200 700 2.13 0.40 4.0 5000900 26 260 30 1.7 760 RR Eri 3300 800 1.84 0.04 3.0 4700900 20 20 3.2 6.0 680 RT Cnc 3300 1350 1.92 0.27 4.0 2100300 4.2 42 5.5 3.0 230 RU Her∗ 2700 700 1.98 0.20 3.4 200005000 50 130 72 13 1230 RW Boo 3200 600 1.90 0.06 4.0 2200300 6.6 66 3.5 2.1 1670 RZ UMa 3100 1250 1.72 0.20 3.4 51001000 7.4 19 21 17 70 ST UMa 3400 1150 1.86 0.09 3.6 2900500 7.5 30 4.0 – – SV Peg 2900 950 1.72 0.20 3.0 100002000 21 21 41 45 330 T Dra 2700 1200 2.00 4.5 4.8 130003000 19 1200 1200 37 780 TU And 3200 1500 1.84 0.35 4.4 75001700 5.9 150 127 – – UY Leo 3300 1350 1.80 0.20 4.2 2000300 4.0 64 23 – – V Eri 3000 1200 1.74 0.45 3.2 51001000 8.1 13 26 65 90 VY UMa 3100 1450 2.22 0.40 4.0 4700900 6.1 24 0.15 0.050 40000 VY UMa∗∗ – – – – – – – – 5.5 – – W Peg 3100 950 2.03 0.35 4.0 53001000 15 150 26 4.1 1100 Y CrB 3100 1300 1.76 0.25 3.2 4500800 6.4 64 7.5 8.1 160 Y Gem 3100 1250 1.70 0.03 2.8 85001800 9.2 5.8 1.0 – – Z Cnc 3200 1000 1.6 0.10 2.0 4200600 12 1.2 1.1 11 55 • Notes. Column (1): name of the source, Col. (2): stellar effective temperature, Col (3): dust shell inner radius temperature, Col (4): dust shell density power-law index, Col (5): dust shell radial optical depth at 550 nm (from to ), Col. (6): logarithm of the dust shell outer-to-inner radii ratio, Col. (7): bolometric luminosity, Col. (8): dust shell inner radius, Col. (9): dust shell outer radius, Col. (10): dust shell mass, Col. (11): dust mas-loss rate, and Col. (12): gas-to-dust ratio. ∗ In the case of RU Her, the sensitivity study indicates 2700 K, although we noted that 2600 K fits correctly the SED (see Fig. 1). ∗∗ In the case of VY UMa, the dust mass estimated for the detached shell is included separately.
4.2 Derivation of related parameters
We derived the bolometric luminosity, the inner and outer radii of the CSE (, ), the radial optical depth as a function of wavelength (see also Appendix D), and the dust mass () from the spatial integration of the density distribution (see eqn. of Sahai et al., 2023).
where =(-1)/(1-), and and are, respectively, the radial optical depth of the shell and the dust mass absorption coefficient at a certain wavelength. In the case of a dust grain mixture, DUSTY considers only one type of grain whose properties are averaged from those of the mixture according to their abundances (see Ivezic & Elitzur, 1995, 1997).
We considered 60 m as the reference wavelength, which is long enough with respect to the assumed grain sizes that grain size does not affect significantly the dust mass absorption coefficient (see e.g. Ysard et al., 2018). The mass absorption coefficients were estimated from their respective complex refractive indexes applying the Mie theory (see e.g. Draine & Lee, 1984; Ivezic & Elitzur, 1997, and references therein).
We obtained the values of for each dust composition used in this study. The different mixtures of O-rich dust has 100 cm2 g-1 (with small differences between them) and the mixture of C-rich dust has 90 cm2 g-1. These estimates are in good agreement with common values from the literature (e.g. Fig. 3 from Ysard et al., 2018).
In the case of VY UMa, we estimated the dust shell masses for the two different components: the present-day CSE (point-source) and the detached shell (extended). This separation allows a more direct comparison with the rest of the sample. The present-day CSE has the lowest dust shell mass among the dusty sources (1.5 ), indicating that the mass-loss process was resumed recently.
The sample presents a wide range of masses, ranging from 1.5 to 1.2 . These mass estimates have a strong dependence on (, i.e. linearly when =2), which is a relatively uncertain parameter.
4.3 Dust mass-loss rates and gas-to-dust ratio.
We performed a comparison between the dust properties obtained in this study and those from the molecular gas component estimated by Alonso-Hernández et al. (2024). For this, we assumed that the expansion velocity of the dust () is similar to that of CO and estimated the dust mass-loss rates by dividing the dust masses over the dust expansion times (=):
Even though the dust masses estimated have a large dependence on , which is uncertain in some cases, the derived dust mass-loss rates are much less dependent of because the expansion time compensates for this dependence (exactly when =2) and our estimated values of are close to 2. We therefore estimate the average gas-to-dust ratios as the ratio between the gas and dust mass-loss rates:
where we used the estimates from Alonso-Hernández et al. (2024). Although the dust expansion velocities are expected to be somewhat greater than those of the gas because of the finite drift velocity of the dust grains relative to the gas, this difference is not significant for the dust mass-loss rates estimates and produces only slight differences (see Groenewegen et al., 1998).
We obtained moderate dust mass-loss rates, in the range 0.05-65 . The lowest estimate corresponds to the present-day CSE of VY UMa (5 ), in which the estimation of is not straightforward due to the presence of two components in the dust, whereas only one was detected in CO. Therefore, we assumed that the CO emission and gas mass-loss rate are related to the present-day dust CSE and used it to estimate , which has a high value of 40000. The rest of values in the sample cover the range 50-1700.
4.4 Trends and comparison with larger samples
In this work, we characterised the dust component in CSEs of uvAGBs and estimate the gas-to-dust ratio by comparing the gas and dust mass-loss rates. The next step is to assess whether these parameters might be sensitive to the UV emission, and its underlying mechanism, or are determined by the stellar properties.
For this purpose, we performed a comparison with one independent stellar property and one extrinsic to the star. For the intrinsic stellar property, we used the variability classification and pulsation period () from Samus’ et al. (2017), summarised for this sample in Alonso-Hernández et al. (2024). For the extrinsic parameter, we used , which is associated with binarity (when 0.06, see Sahai et al., 2022) and large UV excesses (see Appendix. D).
Fig. 4 shows the HRD for our sample, in which the dust shell masses and variability classification are also shown. We found that “naked” stars are located at the bottom-left corner (low luminosity and high temperature), which is consistent with an early evolutionary stage and indicates that they are not undergoing a significant mass-loss process yet. On the other hand, AGB stars with larger dust masses are located at the top-right corner (high luminosity and low temperature), these sources are in a more evolved stage and have more mass accumulated in their CSE. It can also be seen an evolutionary trend over the AGB, where irregular pulsators are located on the bottom-left corner, Miras pulsators are mostly located at the top-right, and Semi-regulars are spread over the other two groups.
Fig. 5 shows the comparisons between the most relevant parameters estimated from the SED modelling with and (defined as corrected from dust attenuation, see appendix D). We evaluate these comparisons qualitatively because the large dependence between these parameters and different methodologies employed in the literature prevents the derivation of simple relationships to compare with (see discussion in Sect. 2 of Höfner & Olofsson, 2018)
We found a decreasing (non-linear) trend of with (see Fig. 5 top-left). This relationship is a trend through the AGB, in which the stellar effective temperatures decrease while pulsation periods increase as the stars climb the AGB.
No clear correlation was found between and (see Fig. 5 middle-top-left), indicating that does not depend directly on the pulsations and related stellar parameters. Low values indicate a lack of warm dust, which might result from a recent decrease in the mass-loss rate or from the accumulation of dust in structures at a certain distance from the star.
We also noted a dichotomy in the values of (see Fig. 5 middle-bottom-left). Miras present standard values with 2.0 on average, whereas semi-regulars systematically present lower values, with 1.8 on average. Even though this result does not have a direct interpretation, the systematically lower values can be related with deviations from homogeneous outflows, in which pulsation driven mass-loss would have more important contribution in semi-regulars than in Miras.
We also found a positive (non-linear) correlation between and (see Fig. 5 bottom-left). The optical depth is proportional to the mass-loss rate, which increases with (see e.g. McDonald & Zijlstra, 2016), and indicates that the dust production rates are higher for more evolved stars.
[b] Source (K) L() AT Dra 3400 2500400 BC CMi 3400 82090 BD Cam 3500 2800600 BY Boo 3400 2100300 CG UMa 3400 2300400 DF Leo 3500 1300200 ome Vir 3400 2300500 RR UMi 3400 1400300 • Notes. Column (1): name of the source, Column (2): stellar effective temperature, Column (3): bolometric luminosity.
On the other hand, no clear correlations were found between the parameters estimated from the SED modelling and (see Fig. 5 right panels). Only a slight decreasing trend of the stellar temperature with might be seen. This correlation, although weak, is expected due in part to the contribution of the stellar photospheric emission in the NUV band increases with , whereas its effect in the FUV band is negligible as it is dominated by accretion or chromospheric emission. Therefore, stars with larger present systematically lower .
Fig. 6 shows the comparison between the gas and dust mass-loss rates estimates for our sample of uvAGBs and those estimated by Wallström et al. (2025) for the Nearby Evolved Stars Survey (NESS, Scicluna et al., 2022) sample, which includes 485 AGB stars. Our CO detected uvAGBs are located at moderate gas and dust mass-loss rates (6 3 and 5 7) and, apart from VY UMa, show a good agreement with values derived from the NESS sample.
We found that our sources are in good agreement with those of NESS and are spread around the trend shown in Wallström et al. (2025). Furthermore, we noted that the two C-rich AGB stars and the four O-rich Miras in our sample are systematically above this trend. We expand the comparison with the NESS sample in Appendix F.
Finally, we searched for correlations between the gas-to-dust ratio () and different parameters (see Fig. 7). We use the linear correlation coefficient () as a simple statistical indicator to identify first-order trends (we refer to 0.7 and 0.70.3 as strong and weak correlations respectively).
We checked the positive correlation between and the ratio between CO intensity and IRAS 60 m flux (=0.37). When rejecting R UMa, which has a significant dispersion in the FIR photometry (see Sect. 3.2), this correlation becomes stronger (=0.63). We also found weak positive correlations of with (=0.51) and (=0.64), whereas a weak anti-correlation was found with (=-0.44).
5 Discussion
In this study, we have characterised the dusty envelopes in a sample of 29 UV emitting AGB stars, which are binary candidates. In particular, we searched for trends between the parameters obtained from SED modelling, the pulsation properties, and the UV emission. We also performed a comparison with the NESS sample of AGB stars, to discern whether uvAGBs present systematic differences in their mass-loss rates or gas-to-dust ratios.
We found that UV emitting AGB stars cannot be clearly differentiated from the rest of AGB stars, on the basis of the their SED shapes, mass-loss rates or gas-to-dust ratios. This result is rather surprising considering the variety of binary-related physical phenomena that may affect the mass-loss, dust formation and enrichment, and gas-to-dust ratios in uvAGBs. In particular:
1. In uvAGBs, the internal UV/X-ray emission may photodissociate the abundance of CO in the innermost regions on the envelope. These abundance depletions can subtly modify the CO line profiles and intensities, remaining unperceived and affecting the gas mass-loss rates estimates (e.g. a factor 2 for T Dra, see Alonso-Hernández et al., 2025).
2. Considering that uvAGBs are likely binary systems, they can suffer alterations in the dust formation process, especially within the wakes of the orbiting close stellar companions, where dust has been proposed to form abundantly (Danilovich et al., 2025). Similarly, the impact of UV emission on the abundances of dust precursor molecular species (e.g. SiO and C2H2, see Van de Sande & Millar, 2019, 2022) may affect dust formation.
3. The presence of binary-induced structures (e.g. spirals and discs), which usually cannot be directly identified by CO line profiles or SED modelling, may lead to incorrect estimates of the mass-loss rates for both dust (see Wiegert et al., 2020) and gas (see Vermeulen et al., 2025) when assuming spherical geometry.
The combination of these phenomena can potentially produce variations in the mass-loss rates of even orders of magnitude. Even though it might be expected that uvAGBs have lower gas-to-dust ratios as suggested by Alonso-Hernández et al. (2024), these phenomena may each other in the overall result. On the other hand, some properties of dust grains (e.g. composition, size) in these systems may differ due to the UV irradiation or effects of stellar companions on the environment.
We also remark that binarity is expected to be common in AGB stars. In particular, the fraction of uvAGBs is at least 10% of the total population (see Alonso-Hernández et al., 2024). Therefore, it is likely that the uvAGBs present in the NESS sample (see Appendix D) are producing an overlap when comparing the samples, avoiding a clear identification of differences.
We also show that modelling dust attenuation at UV wavelengths is necessary for extracting intrinsic UV emission properties (see appendix. D). Optically thick CSEs can prevent even the detection of sources with intrinsically intense UV emission, resulting in an underestimate in the percentages of uvAGBs (as suggested by Montez et al., 2017).
We also highlight the effect of dust attenuation on the 0.06 threshold proposed by Sahai et al. (2022) as a criteria to discern cases in which the UV excesses are clearly related with accretion processes. The observed may well differ from the intrinsic ratio due to the difference in the optical depths between both bands. The optical depth in the FUV band is larger than in the NUV band, leading to intrinsic larger than observed. O-rich AGB stars present a larger difference in the optical depth between both bands (see appendix. D).
The wavelength dependence of dust attenuation also affects the overall shape of the UV spectra, affecting both continuum measurements and line intensity ratios. Although this effect is small when comparing spectral lines to their adjoining continuum emission (e.g. Guerrero & Ortiz, 2020), it can be important when measuring the continuum over a broad spectral range or when estimating ratios between distant spectral lines due to the difference in attenuation (e.g. Ortiz et al., 2019).
We found a weak anticorrelation between the gas-to-dust ratio and , although we acknowledge that the uncertainties associated with the dust attenuation correction and the intrinsic variability of the UV emission in these sources may well be hiding a stronger anticorrelation with . Furthermore, the correlations between the gas-to-dust ratio with the mass-loss rate and stellar pulsations are stronger (see Fig. 7).
In this context, we remark that uvAGBs are predominantly semi-regulars (see Alonso-Hernández et al., 2024) and have moderate or low mass-loss rates (the dust attenuation produces an observational bias to detect UV emission in AGB stars with low opacities), which have systematically lower gas-to-dust ratios. Future studies are required to better understand the effect of UV emission on gas-to-dust ratios.
6 Summary & conclusions
In this paper we studied the dust properties of UV emitting AGB stars estimated from SED modelling. We included the reduction of Herschel/PACS images and performed aperture photometry, which was useful to check the angular sizes and density distributions of the CSEs and to model the extended detached shell of the C-rich AGB star VY UMa.
We obtained moderate dust mass-loss rates in the range – and gas-to-dust ratios ranging from 50 to 1700 (excluding VY UMa, see Table 1), which are intermediate values in comparison with those estimated for the NESS sample (see Fig. 4.3 and Appendix F). Our sample of uvAGBs presents similar mass-loss rates and shows, at least qualitatively, the same trends as larger samples of AGB stars. However, it is expected that larger samples also contain some uvAGBs as they are at least 10% of the population(see Alonso-Hernández et al., 2024), which may prevent us from identifying clear differences.
We found a weak anticorrelation (=-0.44) between the gas-to-dust ratio and . This result indicates that the gas-to-dust ratio of uvAGBs with larger might be more affected by UV emission and/or by the presence of a stellar companions. However, we note that is affected by the uncertainties on the dust attenuation correction as well as the intrinsic UV variability found in uvAGBs. On the other hand, the gas-to-dust ratio is also affected by pulsations and mass-loss rates. Therefore, the relationship between the gas-to-dust ratio and the UV emission is not simple.
It is important to obtain not only samples with confirmed stellar companions, but also of single (or not interacting) stellar systems, to better understand the effect of binarity on mass-loss processes. Future observations with high angular resolution are required to explore the dust forming regions surrounding these stars, identify the presence of stellar companions, and constrain their effect on dust formation and enrichment.
Infrared interferometry is useful to observe the dust distribution in the stellar vicinity, which is key to identifying whether dust formation can be correctly described by radiatively driven winds (Răstău et al., 2026), or to indicate the presence of stellar companions (see e.g. Planquart et al., 2024). Sub-mm interferometry has already shown large capabilities to identify binary-induced structure and modified dynamic in the interior of CSEs (see e.g. Decin et al., 2020). Moreover, larger spectral coverage is required for robust SED modelling, especially extending into the far-infrared, which traces the outermost regions of the CSEs where a significant mass of cool dust may reside.
We also described the effects of dust attenuation on the observed UV emission and the importance of correct modelling when analysing photometric or spectroscopic observations. Correct modelling of dust attenuation can provide a better understanding of the nature of the UV source (e.g. UV excesses and ) as it is not negligible in some cases. Furthermore, dust attenuation can prevent the detection of uvAGBs.
The detectability of UV emission is biased towards AGB stars with low mass-loss rates (and low opacities), which have systematically lower gas-to-dust ratios. A systematic study, using a larger sample, in which binary and single AGB stars are identified, and covering wider ranges of parameters, would help to refine the observed trends with the mass-loss properties.
Data availability
The table containing the SED photometry will be available at the CDS.
Acknowledgements.
We thank the referee, Roberto Ortiz, whose valuable comments helped us improve the quality of the manuscript. This work is part of the I+D+i projects PID2019-105203GB-C22, PID2022-137241NB-C42 and PID2023-146056NB-C22 funded by Spanish MCIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe”. J.A.H. is supported by INTA grant PRE_MDM_05 and acknowledges CSIC grant iMOVE 23023. R.S.’s contribution to the research described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA, and funded in part by NASA via various ROSES/ADAP awards and HST/GO awards (administered by STScI). Herschel is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA. ISO is an ESA project with instruments funded by ESA Member States and with the participation of ISAS and NASA. Some of the data presented in this paper were obtained from the Multimission Archive at the Space Telescope Science Institute (MAST). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555. Support for MAST for non-HST data is provided by the NASA Office of Space Science via grant NAG5-7584 and by other grants and contracts.References
- Agúndez et al. (2020) Agúndez, M., Martínez, J. I., de Andres, P. L., Cernicharo, J., & Martín-Gago, J. A. 2020, A&A, 637, A59
- Alonso-Hernández et al. (2025) Alonso-Hernández, J., Sánchez Contreras, C., Agúndez, M., et al. 2025, A&A, 698, A319
- Alonso-Hernández et al. (2024) Alonso-Hernández, J., Sánchez Contreras, C., & Sahai, R. 2024, A&A, 684, A77
- Aringer et al. (2016) Aringer, B., Girardi, L., Nowotny, W., Marigo, P., & Bressan, A. 2016, MNRAS, 457, 3611
- Aringer et al. (2019) Aringer, B., Marigo, P., Nowotny, W., et al. 2019, MNRAS, 487, 2133
- Balick & Frank (2002) Balick, B. & Frank, A. 2002, ARA&A, 40, 439
- Begemann et al. (1997) Begemann, B., Dorschner, J., Henning, T., et al. 1997, ApJ, 476, 199
- Castro-Carrizo et al. (2010) Castro-Carrizo, A., Quintana-Lacaci, G., Neri, R., et al. 2010, A&A, 523, A59
- Clegg et al. (1996) Clegg, P. E., Ade, P. A. R., Armand, C., et al. 1996, A&A, 315, L38
- Cox et al. (2012) Cox, N. L. J., Kerschbaum, F., van Marle, A. J., et al. 2012, A&A, 537, A35
- Danilovich et al. (2025) Danilovich, T., Samaratunge, N., Mori, Y. L., et al. 2025, A&A, 704, A341
- De Angeli et al. (2023) De Angeli, F., Weiler, M., Montegriffo, P., et al. 2023, A&A, 674, A2
- de Graauw et al. (1996) de Graauw, T., Haser, L. N., Beintema, D. A., et al. 1996, A&A, 315, L49
- De Marco (2009) De Marco, O. 2009, PASP, 121, 316
- Decin et al. (2020) Decin, L., Montargès, M., Richards, A. M. S., et al. 2020, Science, 369, 1497
- Dehaes et al. (2007) Dehaes, S., Groenewegen, M. A. T., Decin, L., et al. 2007, MNRAS, 377, 931
- Doi et al. (2015) Doi, Y., Takita, S., Ootsubo, T., et al. 2015, PASJ, 67, 50
- Draine & Lee (1984) Draine, B. T. & Lee, H. M. 1984, ApJ, 285, 89
- Duquennoy & Mayor (1991) Duquennoy, A. & Mayor, M. 1991, A&A, 248, 485
- Ferrarotti & Gail (2006) Ferrarotti, A. S. & Gail, H. P. 2006, A&A, 447, 553
- Gordon et al. (2023) Gordon, K. D., Clayton, G. C., Decleir, M., et al. 2023, ApJ, 950, 86
- Graciá-Carpio et al. (2015) Graciá-Carpio, J., Wetzstein, M., & Roussel, H. 2015, arXiv e-prints, arXiv:1512.03252
- Groenewegen (2022) Groenewegen, M. A. T. 2022, A&A, 659, A145
- Groenewegen et al. (1998) Groenewegen, M. A. T., Whitelock, P. A., Smith, C. H., & Kerschbaum, F. 1998, MNRAS, 293, 18
- Gry et al. (2003) Gry, C., Swinyard, B., Harwood, A., et al. 2003, The ISO Handbook, Volume III - LWS - The Long Wavelength Spectrometer
- Guerrero & Ortiz (2020) Guerrero, M. A. & Ortiz, R. 2020, MNRAS, 491, 680
- Hanner (1988) Hanner, M. 1988, Grain optical properties., In NASA, Washington, Infrared Observations of Comets Halley and Wilson and Properties of the Grains p 22-49 (SEE N89-13330 04-89)
- Henning et al. (1995) Henning, T., Begemann, B., Mutschke, H., & Dorschner, J. 1995, A&AS, 112, 143
- Heras & Hony (2005) Heras, A. M. & Hony, S. 2005, A&A, 439, 171
- Höfner & Olofsson (2018) Höfner, S. & Olofsson, H. 2018, A&A Rev., 26, 1
- Ivezic & Elitzur (1995) Ivezic, Z. & Elitzur, M. 1995, ApJ, 445, 415
- Ivezic & Elitzur (1997) Ivezic, Z. & Elitzur, M. 1997, MNRAS, 287, 799
- Ivezic et al. (1999) Ivezic, Z., Nenkova, M., & Elitzur, M. 1999, arXiv e-prints, astro
- Jenness et al. (2002) Jenness, T., Stevens, J. A., Archibald, E. N., et al. 2002, MNRAS, 336, 14
- Jones et al. (2014) Jones, O. C., Kemper, F., Srinivasan, S., et al. 2014, MNRAS, 440, 631
- Kemper et al. (2002) Kemper, F., de Koter, A., Waters, L. B. F. M., Bouwman, J., & Tielens, A. G. G. M. 2002, A&A, 384, 585
- Kessler et al. (2003) Kessler, M. F., Mueller, T. G., Leech, K., et al. 2003, The ISO Handbook, Volume I - Mission & Satellite Overview
- Kessler et al. (1996) Kessler, M. F., Steinz, J. A., Anderegg, M. E., et al. 1996, A&A, 315, L27
- Kwitter & Henry (2022) Kwitter, K. B. & Henry, R. B. C. 2022, PASP, 134, 022001
- Leech et al. (2003) Leech, K., Kester, D., Shipman, R., et al. 2003, The ISO Handbook, Volume V - SWS - The Short Wavelength Spectrometer
- Martin et al. (2005) Martin, D. C., Fanson, J., Schiminovich, D., et al. 2005, ApJ, 619, L1
- Mathis et al. (1977) Mathis, J. S., Rumpl, W., & Nordsieck, K. H. 1977, ApJ, 217, 425
- McDonald & Zijlstra (2016) McDonald, I. & Zijlstra, A. A. 2016, ApJ, 823, L38
- Mečina et al. (2020) Mečina, M., Aringer, B., Nowotny, W., et al. 2020, A&A, 644, A66
- Mečina et al. (2014) Mečina, M., Kerschbaum, F., Groenewegen, M. A. T., et al. 2014, A&A, 566, A69
- Miszalski et al. (2009) Miszalski, B., Acker, A., Moffat, A. F. J., Parker, Q. A., & Udalski, A. 2009, A&A, 496, 813
- Montez et al. (2017) Montez, Rodolfo, J., Ramstedt, S., Kastner, J. H., Vlemmings, W., & Sanchez, E. 2017, ApJ, 841, 33
- Neugebauer et al. (1984) Neugebauer, G., Habing, H. J., van Duinen, R., et al. 1984, ApJ, 278, L1
- Nordhaus & Blackman (2006) Nordhaus, J. & Blackman, E. G. 2006, MNRAS, 370, 2004
- Ochsenbein et al. (2000) Ochsenbein, F., Bauer, P., & Marcout, J. 2000, A&AS, 143, 23
- Onaka et al. (1999) Onaka, T., de Jong, T., Yamamura, I., Cami, J., & Tanab’e, T. 1999, in ESA Special Publication, Vol. 427, The Universe as Seen by ISO, ed. P. Cox & M. Kessler, 381
- Ortiz & Guerrero (2016) Ortiz, R. & Guerrero, M. A. 2016, MNRAS, 461, 3036
- Ortiz & Guerrero (2021) Ortiz, R. & Guerrero, M. A. 2021, ApJ, 912, 93
- Ortiz et al. (2019) Ortiz, R., Guerrero, M. A., & Costa, R. D. D. 2019, MNRAS, 482, 4697
- Ott (2010) Ott, S. 2010, in Astronomical Society of the Pacific Conference Series, Vol. 434, Astronomical Data Analysis Software and Systems XIX, ed. Y. Mizumoto, K. I. Morita, & M. Ohishi, 139
- Pegourie (1988) Pegourie, B. 1988, A&A, 194, 335
- Piazzo et al. (2015) Piazzo, L., Calzoletti, L., Faustini, F., et al. 2015, MNRAS, 447, 1471
- Pilbratt et al. (2010) Pilbratt, G. L., Riedinger, J. R., Passvogel, T., et al. 2010, A&A, 518, L1
- Planquart et al. (2024) Planquart, L., Paladini, C., Jorissen, A., et al. 2024, A&A, 687, A306
- Poglitsch et al. (2010) Poglitsch, A., Waelkens, C., Geis, N., et al. 2010, A&A, 518, L2
- Ramstedt et al. (2012) Ramstedt, S., Montez, R., Kastner, J., & Vlemmings, W. H. T. 2012, A&A, 543, A147
- Rodrigo et al. (2024) Rodrigo, C., Cruz, P., Aguilar, J. F., et al. 2024, A&A, 689, A93
- Rodrigo & Solano (2020) Rodrigo, C. & Solano, E. 2020, in XIV.0 Scientific Meeting (virtual) of the Spanish Astronomical Society, 182
- Rodrigo et al. (2012) Rodrigo, C., Solano, E., & Bayo, A. 2012, SVO Filter Profile Service Version 1.0, IVOA Working Draft 15 October 2012
- Răstău et al. (2026) Răstău, V., Paladini, C., Drevon, J., et al. 2026, A&A, 705, A127
- Sahai et al. (2023) Sahai, R., Bujarrabal, V., Quintana-Lacaci, G., et al. 2023, ApJ, 943, 110
- Sahai et al. (2008) Sahai, R., Findeisen, K., Gil de Paz, A., & Sánchez Contreras, C. 2008, ApJ, 689, 1274
- Sahai et al. (2007) Sahai, R., Morris, M., Sánchez Contreras, C., & Claussen, M. 2007, AJ, 134, 2200
- Sahai et al. (2011a) Sahai, R., Morris, M. R., & Villar, G. G. 2011a, AJ, 141, 134
- Sahai et al. (2011b) Sahai, R., Neill, J. D., Gil de Paz, A., & Sánchez Contreras, C. 2011b, ApJ, 740, L39
- Sahai et al. (2018) Sahai, R., Sánchez Contreras, C., Mangan, A. S., et al. 2018, ApJ, 860, 105
- Sahai et al. (2022) Sahai, R., Sanz-Forcada, J., Guerrero, M., Ortiz, R., & Contreras, C. S. 2022, Galaxies, 10, 62
- Sahai et al. (2016) Sahai, R., Sanz-Forcada, J., & Sánchez Contreras, C. 2016, in Journal of Physics Conference Series, Vol. 728, Journal of Physics Conference Series, 042003
- Sahai et al. (2015) Sahai, R., Sanz-Forcada, J., Sánchez Contreras, C., & Stute, M. 2015, ApJ, 810, 77
- Sahai & Trauger (1998) Sahai, R. & Trauger, J. T. 1998, AJ, 116, 1357
- Samus’ et al. (2017) Samus’, N. N., Kazarovets, E. V., Durlevich, O. V., Kireeva, N. N., & Pastukhova, E. N. 2017, Astronomy Reports, 61, 80
- Scicluna et al. (2022) Scicluna, P., Kemper, F., McDonald, I., et al. 2022, MNRAS, 512, 1091
- Sloan et al. (2003a) Sloan, G. C., Kraemer, K. E., Goebel, J. H., & Price, S. D. 2003a, ApJ, 594, 483
- Sloan et al. (2003b) Sloan, G. C., Kraemer, K. E., Price, S. D., & Shipman, R. F. 2003b, ApJS, 147, 379
- Sloan et al. (2025) Sloan, G. C., Kraemer, K. E., & Volk, K. 2025, ApJS, 279, 15
- Sloan & Price (1995) Sloan, G. C. & Price, S. D. 1995, ApJ, 451, 758
- Speck et al. (2000) Speck, A. K., Barlow, M. J., Sylvester, R. J., & Hofmeister, A. M. 2000, A&AS, 146, 437
- Stanghellini et al. (2016) Stanghellini, L., Shaw, R. A., & Villaver, E. 2016, ApJ, 830, 33
- Suh (2018) Suh, K.-W. 2018, Journal of Korean Astronomical Society, 51, 155
- Tachibana et al. (2023) Tachibana, K., Miyata, T., Kamizuka, T., et al. 2023, PASJ, 75, 489
- Tielens (2005) Tielens, A. G. G. M. 2005, The Physics and Chemistry of the Interstellar Medium
- Ueta et al. (2000) Ueta, T., Meixner, M., & Bobrowsky, M. 2000, ApJ, 528, 861
- Van de Sande & Millar (2019) Van de Sande, M. & Millar, T. J. 2019, ApJ, 873, 36
- Van de Sande & Millar (2022) Van de Sande, M. & Millar, T. J. 2022, MNRAS, 510, 1204
- van Marle et al. (2014) van Marle, A. J., Cox, N. L. J., & Decin, L. 2014, A&A, 570, A131
- Vermeulen et al. (2025) Vermeulen, O., Esseldeurs, M., Malfait, J., et al. 2025, A&A, 700, A85
- Wallström et al. (2025) Wallström, S. H. J., Scicluna, P., Srinivasan, S., et al. 2025, A&A, 704, A276
- Waters (2011) Waters, L. B. F. M. 2011, in Astronomical Society of the Pacific Conference Series, Vol. 445, Why Galaxies Care about AGB Stars II: Shining Examples and Common Inhabitants, ed. F. Kerschbaum, T. Lebzelter, & R. F. Wing, 227
- Wiegert et al. (2020) Wiegert, J., Groenewegen, M. A. T., Jorissen, A., Decin, L., & Danilovich, T. 2020, A&A, 642, A142
- Ysard et al. (2018) Ysard, N., Jones, A. P., Demyk, K., Boutéraon, T., & Koehler, M. 2018, A&A, 617, A124
Appendix A Semi-extended sources
Fig. 8 shows radial surface brightness profiles for the four semi-extended sources presented in this study (RU Her, SV Peg, T Dra and V Eri). In addition, we created DUSTY synthetic images for their best-fit models in the effective wavelength of each Herschel/PACS filter and convolved them with their respective PSFs. Then, we radially averaged the surface brightness, in both the observed and the synthetic images, in radial annuli of 1″and compared them. The uncertainties were estimated as the sum of the standard deviation of the flux and the average instrumental error in each annulus.
The best-fit DUSTY models (see Sect. 4.1) reproduce correctly the angular sizes after convolving with the correspondent PSF. In the four cases the difference between the angular sizes and the Herschel/PACS PSFs can be appreciated better on the B filter (70 m) due to its higher angular resolution.
Appendix B VY UMa detached shell
We present a more detailed analysis of the SED modelling for the particular case of VY UMa, which has a large-scale detached shell in addition to a compact, bright central emission component that represents the present-day mass-loss rate. This extended emission was previously identified in the Herschel/PACS images (Cox et al. 2012; van Marle et al. 2014) and manifested in the SED as the large excesses in the far-infrared (see the case of VY UMa in Fig. 1).
The contribution of the detached shell is only noticed in the SED in the far-infrared because it is composed of very cold dust (10–100K), whose emission is concentrated in the 30–300 m region, and it does not modify the SED at shorter wavelengths. Therefore, we used the Herschel/PACS fluxes, which isolates the present-day mass-loss (black filled circles in Fig. 1) and the extended emission from the detached shell (black filled diamonds in Fig. 1), and rejected IRAS and AKARI because they included partially the emission form the detached shell.
To correctly fit the SED, we followed a similar methodology to that presented in Mečina et al. (2014, 2020) to separate the contribution from the present-day mass-loss and the detached shell. We started by fitting the SED of the present-day mass-loss (including Herschel/PACS fluxes and all the photometric observations at ¡ 30 m) following the same procedure as in the rest of the sources. Once the best-fit solution for the present-day mass-loss was found, we modified the normalised density distribution including a peak at 3200–8000, which is equivalent to 31–78″ considering the source distance and in agreement with the location of the detached shell (see Fig. 9).
This model reproduces satisfactorily both the FIR excesses and the Herschel/PACS radial surface brightness profiles. It allows us to estimate an overall density and dust mass of the extended detached shell. However, we note that this single 1D model is not able to reproduce the asymmetrical shape of the detached shell, which would require a more advanced 3D radiative transfer model.
Appendix C Summary of Herschel/PACS observations
We present the compilation of photometric measurements for all the individual Herschel/PACS observation employed in this study. Table 3 shows the technical information and the individual measurements obtained from aperture photometry in the two employed bands of each observation.
[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) AT Dra 1342190578 55239.274 152 chop 1.190.03 – 0.2080.009 1342191845 55264.457 152 chop 1.230.03 – 0.2580.008 EY Hya 1342196885 55338.979 26960 parallel – 2.510.29 0.940.09 1342196886 55339.293 24690 parallel – 2.510.29 0.940.09 1342197282 55345.144 26960 parallel – 1.820.21 0.750.10 1342197283 55345.457 24690 parallel – 1.820.21 0.750.10 1342253344 56211.138 52 scan map 4.050.01 – 0.730.03 1342253500 56215.995 52 scan map 4.140.02 – 0.970.17 IN Hya 1342197307 55345.979 26960 parallel – 0.250.06 0.13 1342197308 55346.292 26960 parallel – 0.250.06 0.13 ome Vir 1342187362 55165.745 161 chop-nod 1.510.03 – 0.260.06 1342198902 55368.940 152 chop-nod 1.480.03 – 0.290.06 1342223317 55741.738 152 chop-nod 1.480.02 – 0.330.05 1342223625 55745.669 152 chop-nod 1.450.02 – 0.270.05 1342223690 55746.726 152 chop-nod 1.510.02 – 0.310.04 1342223885 55752.614 152 chop-nod 1.510.03 – 0.260.05 1342223925 55753.611 152 chop-nod 1.580.03 – 0.260.06 R LMi 1342254139 56230.736 52 scan map 11.720.02 – 2.450.05 R UMa 1342221940 55712.986 52 scan map 2.810.05 – 0.470.05 1342246162 56073.792 52 scan map 2.990.04 – 0.560.04 1342246187 56074.789 52 scan map 3.120.06 – 0.590.05 1342246747 56085.991 52 scan map 3.270.06 – 0.690.04 1342246935 56088.951 52 scan map 3.330.06 – 0.570.02 RR Eri 1342190602 55239.380 152 chop-nod 1.750.02 – 0.350.05 RR UMi 1342254046 56229.690 52 scan map 2.970.01 – 0.530.02 RT Cnc 1342253343 56211.136 52 scan map 2.300.07 – 0.430.04 1342253469 56215.813 52 scan map 2.440.06 – 0.560.06 RU Her 1342190522 55239.065 152 chop-nod 8.370.03 – 1.430.03 1342191825 55264.384 152 chop-nod 8.220.03 – 1.380.04 1342228878 55822.857 152 chop-nod 6.610.03 – 1.140.03 ST UMa 1342247296 56101.709 52 scan map 1.630.03 – 0.320.09 1342247425 56103.683 52 scan map 1.630.04 – 0.190.06 SV Peg 1342220943 55696.010 152 chop-nod 13.030.05 – 1.870.16 1342221173 55697.791 152 chop-nod 13.170.04 – 1.830.12 1342222726 55729.050 152 chop-nod 12.880.05 – 1.840.11 • Notes. Column (1): Source, Col. (2): Obs ID, Col. (3): OBs date, Col. (4): exposure duration, Col. (5): observational mode, Col. (6): Flux in band B, Col. (7): Flux in band G, Col. (8): Flux in band R. Herschel/PACS B, G and R filters have effective wavelengths of 72, 103 and 167 m, and effective spectral widths of 21, 31, and 70 m respectively.
[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) T Dra 1342187357 55165.608 161 chop-nod 8.470.02 – 1.660.03 1342190576 55239.267 152 chop-nod 7.760.02 – 1.550.03 1342222753 55729.353 152 chop-nod 8.660.02 – 1.620.03 1342232432 55882.643 152 chop-nod 10.000.02 – 2.080.03 1342232465 55883.675 152 chop-nod 11.100.02 – 2.020.03 1342233065 55891.711 152 chop-nod 11.130.02 – 2.020.03 1342233205 55892.709 152 chop-nod 11.020.02 – 2.030.03 1342233356 55896.707 152 chop-nod 11.040.02 – 1.990.03 1342233587 55899.624 152 chop-nod 11.070.02 – 1.980.03 1342234216 55900.630 152 chop-nod 10.930.02 – 2.000.03 1342240701 55992.767 152 chop-nod 9.090.02 – 1.690.03 1342241298 55999.595 152 chop-nod 8.930.02 – 1740.03 V Eri 1342189943 55227.184 283 scan map 12.610.04 – 1.900.16 1342189944 55227.188 283 scan map 12.610.04 – 1.900.16 VY UMa 1342210606 55530.947 881 scan map 1.200.04 – 0.190.02 1342210607 55530.958 881 scan map 1.200.04 – 0.190.02 1342243219 56008.669 733 scan map – 0.640.02 0.230.03 1342243220 56008.678 733 scan map – 0.640.02 0.230.03 W Peg 1342222729 55729.060 152 chop-nod 6.610.03 – 1.280.04 Z Cnc 1342251911 56200.817 52 scan map 1.330.01 – 0.200.13
Appendix D UV excesses and dust attenuation
We computed synthetic stellar photometry on the GALEX bands and compared them with the observed fluxes to estimate the UV excesses (similarly to Sahai et al. 2008; Ortiz & Guerrero 2016). We noted that GALEX NUV and FUV photometric bands fall outside the wavelength range covered by the COMARCS models used in this work (see Sect. 3). Therefore, the stellar fluxes were estimated by extrapolating the input spectra following a blackbody distribution and convolving it with the transmission curves of their respective filters. We acknowledge that the real stellar spectrum differs from a blackbody approximation mainly due to the presence of (i) molecular absorption bands, and (ii) emission-line features due to possible chromospheric emission, that can produce a significant effect on the shape of the stellar spectra in the UV. The features in (ii) are expected to have a much stronger effect in the FUV (compared to the NUV) where the photospheric contribution is relatively much weaker. Note that the sole purpose of this simple approximation is to explore whether is a reasonable proxy for the FUV excess.
In addition, the UV spectral range is specially affected by circumstellar dust attenuation, which might imply a large difference between the intrinsic and the observed UV emission, therefore, affecting the estimation of the UV excesses. Moreover, the wavelength dependency of the opacity produces a variation of the .
Fig. 11 shows the opacity curves for the generic dust composition of C-rich AGB stars and O-rich AGB stars described in Sect. 3.1 provided by DUSTY as well as the ISM from Gordon et al. (2023). In O-rich AGB stars the dust opacity is dominated by silicates and oxides, resulting in . In C-rich AGB stars the dust opacity is dominated by amorphous carbon, which shows the 2200Å NUV bump, resulting in . Finally, the opacity curve of the ISM also displays the NUV bump, although with a different overall shape, resulting in .
The observed UV fluxes were corrected for the dust attenuation in order to estimate their intrinsic values, assuming that the UV source is located inside the dust envelope and, therefore, the flux is attenuated by the full radial optical depth of the dust envelope. This assumption is reasonable considering that the UV emission is expected to be produced by the presence of an interacting close stellar companion, located at a few stellar radii and inside the dust formation zone.
Fig. 12 shows the dependence of the opacity correction factor with the optical depth and its effect on . The optical depth is larger in the FUV band than in the NUV band, leading to intrinsic larger than observed. Furthermore, the opacity correction factor increases exponentially with the optical depth. The effect on O-rich AGB stars is significant for large opacities, whereas in C-rich AGB stars the effect remains very small in this opacity range, due to the differences between the opacity on NUV and FUV bands.
Fig 13 shows the comparison between and the UV excesses. It also shows linear fits between these parameters. There is a weak correlation (=0.39) between and the NUV excess, which becomes stronger (=0.62) after correcting for the dust attenuation. In contrast, the correlation between and the FUV excess is strong (=0.85) and becomes even stronger (=0.89) after the correction. This correlation shows that is a good tracer of UV excesses and can be used independently to explore trends.
This result is expected because small UV excesses indicate a larger contribution of the stellar component to , and stellar photospheric emission is characterised by low . As described in Sect. 1, both UV excesses and are related to the nature of the UV emission: small UV excesses and indicate an intrinsic origin (i.e. chromospheric activity), whereas large UV excesses and indicate an extrinsic origin (i.e. accretion activity). We highlight that these correlations can be seen despite the scatter produced by UV emission variability, and they become more pronounced when corrected for dust attenuation.
[b] Source NUV FUV NUV FUV NUV FUV (Obs.) (Obs.) (SED) (SED) (excess) (excess) (Obs.) (Jy) (Jy) (Jy) (Jy) () AT Dra 2190 177 1700 7.6e 1.29 2.30 0.080 – – 0.080 BC CMi(∗) 1013 55 720 2.9 1.41 1.90 0.054 – – 0.054 BD Cam(∗) 5970 2330 4200 3.2 1.42 7.28 0.390 – – 0.390 BY Boo 3140 214 2100 8.3 1.50 2.58 0.068 – – 0.068 CG UMa 1498 108 1200 4.9 1.25 2.20 0.072 – – 0.072 DF Leo 590 36 1000 7.6 0.59 0.47 0.061 – – 0.061 EY Hya 110 82 23 2.8 4.8 290 0.740 0.99 1.67 1.461 FH Vir 400 49 370 1.4 1.08 3.5 0.120 0.14 0.24 0.133 IN Hya 620 44 1200 8.6 0.52 0.51 0.071 0.11 0.19 0.077 ome Vir(∗) 2950 179 2100 9.3 1.40 1.92 0.061 – – 0.061 R LMi(b) 40 2.4 7.6 5.2 5 46 0.040 4.50 7.63 0.9 R UMa 90 50 30 5.3 3.0 94 0.6 1.13 1.91 1.3 RR Eri 430 28 300 8.4 1.4 3.3 0.064 0.08 0.14 0.068 RR UMi 7200 1870 4100 1.8 1.76 1.04 0.26 – – 0.26 RT Cnc 190 32 170 3.7 1.12 9 0.19 0.76 1.29 0.32 RU Her(b) 80 5.75 27 5.8 3.0 10 0.08 0.56 0.95 0.12 RW Boo 410 23 180 5.3 2.3 4.3 0.057 0.17 0.29 0.064 RZ UMa(b) 18 8 22 3.6 0.82 22 0.4 0.51 0.86 0.6 ST UMa(∗) 994 168 680 2.7 1.46 6.2 0.17 0.25 0.43 0.20 SV Peg(b) 29 5.75 39 4.6 0.74 13 0.21 0.55 0.91 0.30 T Dra(a,b) 32 7.9 2.6 4.2 12.3 190 0.25 7.76 8.05 0.33 TU And(b) 15 5.75 19 4.3 0.8 13 0.96 0.98 1.67 1.91 UY Leo 150 27 29 7.6 5.2 36 0.18 0.56 0.95 0.27 V Eri 100 67 34 4.0 2.9 170 0.67 1.41 2.39 1.8 VY UMa 140 8 110 3.0 1.3 2.7 0.06 0.69 0.72 0.06 W Peg 180 13 41 5.8 4.4 22 0.065 0.98 1.67 0.13 Y CrB(b) 12 5.75 16 3.3 0.8 17 0.67 0.70 1.19 1.09 Y Gem 6000 5000 4.6 1.0 1300 5000 0.86 0.08 0.14 0.91 Z Cnc 73 12 9.2 2.4 7.9 5.0 0.17 0.28 0.48 0.21 • Notes. Column (1): Source, Col. (2): Averaged observed NUV flux, Col. (3): Averaged observed FUV flux, Col. (4): NUV flux expected from SED modelling, Col. (5): FUV flux expected from SED modelling, Col. (6): NUV excess, Col. (7): FUV excess, Col. (8): Averaged ratio between FUV and NUV fluxes, Col. (9): Optical depth at NUV effective wavelength, Col. (10): Optical depth at FUV effective wavelength, Col. (11): Averaged ratio between FUV and NUV fluxes corrected from dust attenuation.
Appendix E Parameter uncertainties and correlations from the SED modelling
We present a detailed description of the procedure used to estimate the uncertainties of the free parameters used in the SED modelling and identify correlations between them. Single parameter uncertainties were estimated as 68% confidence intervals from the normalised log-likelihood distribution . We defined the likelihood function as:
where are the free parameters. On the other hand, we explore degeneracies between pairs of parameters from their distributions. This method allowed us to identify which parameters are degenerated and visualise the degree to which these degeneracies affect the value.
Fig. 14 shows the corner plot obtained for EY Hya as an example of the estimation of the uncertainties associated with the free parameters from the SED modelling and the identification of the correlations between them. We found the following sets of positive correlations: –, –, –, –, and –. On the other hand, we found the following set of anticorrelations: –, –, and –.
These correlations are quite linear in these ranges of the parameters, although there are two exceptions: (i) log(Y) that presents non-linearities with and and (ii) that presents sharp upper limits, this limit is related with an emission excess in optical wavelengths that cannot be compensated by the rest of parameters. Similar correlations were found for the rest of dusty sources and can be generalised for the modelling procedure.
Appendix F Comparison with NESS sample
We present a complementary comparison with the Nearby Evolved Stars Survey (NESS, Scicluna et al. 2022), which contains 485 nearby AGB stars. We first compared the CO(=2-1) and IRAS 60 m fluxes between our sample and NESS (obtained from Wallström et al. 2025) to check whether uvAGBs have statistically lower CO intensity at similar IRAS 60 m as suggested by Alonso-Hernández et al. (2024) using the CO(=1-0).
Fig. 15 shows the comparison between CO(=2-1) integrated intensity and IRAS 60 m flux. We used the CO(=2-1) intensities as they are available for both our sample and NESS. We scaled NESS CO(=2-1) intensities to match our intensities for the three commonly detected in CO in both samples (R LMi, SV Peg and W Peg).
In the case of unresolved and optically thin envelopes the CO intensity is proportional to the infrared flux (i.e. slope of 1 in logarithmic scale). Nevertheless, we noted that some scattering in this relationship is expected due to optically depth, partially resolved sources and ISM pollution. We found that our uvAGBs with the largest IRAS 60 m fluxes in our sample overlap with the trend described by NESS sources. However, our sources with lowest IRAS 60 m fluxes are systematically under this trend, although they fall outside the IRAS 60 m range covered by NESS. The resulting fits indicate that our sources have indeed lower CO(=2-1) intensities as well as CO(=1-0), as previously inferred by Alonso-Hernández et al. (2024).
We also compared the correlations that we found in Fig. 7 for the gas-to-dust ratio with (i) the ratio between CO and IRAS 60 m and (ii) with . Our estimates are in the upper part of those covered by NESS, although this parameter depends in the methodology to estimate the mass-loss rates and it is sensitive to systematic effects. In the case of , we used the values without dust attenuation correction because we do not have the dust optical depths for the NESS sources.
Fig. 16 shows the comparison between the gas-to-dust ratio and the ratio between CO and IRAS 60 m and . It is shown that correlates with the ratio between CO and infrared emission, although some outliers (likely related with pollution in the IRAS 60 m images) are present. We noted that is not a empirical value and depends on the employed assumptions and methodology. Therefore, it is likely the presence of some systematic differences between our estimates and NESS. On the other hand, anticorrelates with , this correlation, although it presents some scatter, which can be related with UV emission variability, and it is not corrected from dust attenuation.
Finally, we crossmatched the NESS catalogue with GALEX and found that among the 485 AGB stars in NESS, 163 were detected in the NUV band and 95 were also detected in the FUV band. These percentages illustrate that large AGB star samples also contain some uvAGBs.

### 27 Mar 2026 v2

@@
 Therefore, the instrumental uncertainties underestimate the intrinsic variability of AGB stars on the far-infrared (this also can apply to IRAS and AKARI photometry).
-The employed Herschel/PACS observations and their respective estimated fluxes are summarised in appendix C.
+The employed Herschel/PACS observations and their respective estimated fluxes are summarised in electronic format at the CDS.
 3 Analysis: dust emission model
@@
 Finally, the radial optical depth, , at a fiducial wavelength of 550 nm, is used to parametrise the optical depth of the CSE across the entire wavelength range covered in the model.
-We used the optical depths at the central wavelengths of the two GALEX bands in order to estimate the effect of dust attenuation in the observed UV emission (see appendix D).
+We used the optical depths at the central wavelengths of the two GALEX bands in order to estimate the effect of dust attenuation in the observed UV emission (see appendix C).
 3.2 Parameter space and SED fitting
@@
 Moreover, we noted that a few sources (R LMi, R UMa and Z Cnc) show a large difference between the fluxes obtained from Herschel/PACS and those from IRAS and AKARI (see Fig.
-1).
+1 and appendix D).
 We noted that, in these cases, IRAS and AKARI fluxes are systematically higher, indicating pollution from an irregular sky brightness due to their more limited angular resolution.
@@
 4.1 Best-fit solutions and uncertainties
-We find that the SEDs of most of our targets (21/29) show an infrared excess and require a dusty CSE (see Figs.
-1 and table 1); we derive additional parameters for these sources in Sects.
+We find that the SEDs of most of our targets (21/29) show an infrared excess and require a dusty CSE (see Fig.
+1, table 1 and appendix D); we derive additional parameters for these sources in Sects.
 4.2 and 4.3.
@@
 This clustering at low can be expected because uvAGBs with low optical depths have a higher chance of being detected via their UV emission due to a low attenuation (see Sect.
-5 and Appendix D).
+5 and Appendix C).
 We found that, in some cases, is rather uncertain because the wavelength coverage of the observations does not extend to the sufficiently long wavelength that is required to probe the coldest dust located in the outermost regions of the envelopes (see also Heras & Hony, 2005).
@@
 4.2 Derivation of related parameters
-We derived the bolometric luminosity, the inner and outer radii of the CSE (, ), the radial optical depth as a function of wavelength (see also Appendix D), and the dust mass () from the spatial integration of the density distribution (see eqn.
-of Sahai et al., 2023).
+We derived the bolometric luminosity, the inner and outer radii of the CSE (, ), the radial optical depth as a function of wavelength (see also Appendix C), and the dust mass () from the spatial integration of the density distribution (see eqn.
+1 of Sahai et al., 2023).
 where =(-1)/(1-), and and are, respectively, the radial optical depth of the shell and the dust mass absorption coefficient at a certain wavelength.
@@
 (2024).
-For the extrinsic parameter, we used , which is associated with binarity (when 0.06, see Sahai et al., 2022) and large UV excesses (see Appendix.
-D).
+For the extrinsic parameter, we used , which is associated with binarity (when 0.06, see Sahai et al., 2022) and large UV excesses (see Appendix C).
 Fig.
@@
 Fig.
-5 shows the comparisons between the most relevant parameters estimated from the SED modelling with and (defined as corrected from dust attenuation, see appendix D).
+5 shows the comparisons between the most relevant parameters estimated from the SED modelling with and (defined as corrected from dust attenuation, see appendix C).
 We evaluate these comparisons qualitatively because the large dependence between these parameters and different methodologies employed in the literature prevents the derivation of simple relationships to compare with (see discussion in Sect.
@@
 In particular, the fraction of uvAGBs is at least 10% of the total population (see Alonso-Hernández et al., 2024).
-Therefore, it is likely that the uvAGBs present in the NESS sample (see Appendix D) are producing an overlap when comparing the samples, avoiding a clear identification of differences.
-We also show that modelling dust attenuation at UV wavelengths is necessary for extracting intrinsic UV emission properties (see appendix.
-D).
+Therefore, it is likely that the uvAGBs present in the NESS sample (see Appendix C) are producing an overlap when comparing the samples, avoiding a clear identification of differences.
+We also show that modelling dust attenuation at UV wavelengths is necessary for extracting intrinsic UV emission properties (see appendix C).
 Optically thick CSEs can prevent even the detection of sources with intrinsically intense UV emission, resulting in an underestimate in the percentages of uvAGBs (as suggested by Montez et al., 2017).
@@
 The optical depth in the FUV band is larger than in the NUV band, leading to intrinsic larger than observed.
-O-rich AGB stars present a larger difference in the optical depth between both bands (see appendix.
-D).
+O-rich AGB stars present a larger difference in the optical depth between both bands (see appendix C).
 The wavelength dependence of dust attenuation also affects the overall shape of the UV spectra, affecting both continuum measurements and line intensity ratios.
@@
 Future studies are required to better understand the effect of UV emission on gas-to-dust ratios.
-6 Summary & conclusions
+6 Summary and conclusions
 In this paper we studied the dust properties of UV emitting AGB stars estimated from SED modelling.
@@
 Data availability
-The table containing the SED photometry will be available at the CDS.
+Tables containing the employed photometry for the SEDs and the photometric fluxes estimated from the individual Herschel/PACS observations are only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr (130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/.
 Acknowledgements.
@@
 is supported by INTA grant PRE_MDM_05 and acknowledges CSIC grant iMOVE 23023.
-R.S.’s contribution to the research described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA, and funded in part by NASA via various ROSES/ADAP awards and HST/GO awards (administered by STScI).
+R.S.’s contribution to the research described here was carried out at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with NASA (80NM0018D0004), and funded in part by NASA via various ROSES/ADAP awards and HST/GO awards (administered by STScI).
 Herschel is an ESA space observatory with science instruments provided by European-led Principal Investigator consortia and with important participation from NASA.
@@
 However, we note that this single 1D model is not able to reproduce the asymmetrical shape of the detached shell, which would require a more advanced 3D radiative transfer model.
-Appendix C Summary of Herschel/PACS observations
-We present the compilation of photometric measurements for all the individual Herschel/PACS observation employed in this study.
-Table 3 shows the technical information and the individual measurements obtained from aperture photometry in the two employed bands of each observation.
-[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) AT Dra 1342190578 55239.274 152 chop 1.190.03 – 0.2080.009 1342191845 55264.457 152 chop 1.230.03 – 0.2580.008 EY Hya 1342196885 55338.979 26960 parallel – 2.510.29 0.940.09 1342196886 55339.293 24690 parallel – 2.510.29 0.940.09 1342197282 55345.144 26960 parallel – 1.820.21 0.750.10 1342197283 55345.457 24690 parallel – 1.820.21 0.750.10 1342253344 56211.138 52 scan map 4.050.01 – 0.730.03 1342253500 56215.995 52 scan map 4.140.02 – 0.970.17 IN Hya 1342197307 55345.979 26960 parallel – 0.250.06 0.13 1342197308 55346.292 26960 parallel – 0.250.06 0.13 ome Vir 1342187362 55165.745 161 chop-nod 1.510.03 – 0.260.06 1342198902 55368.940 152 chop-nod 1.480.03 – 0.290.06 1342223317 55741.738 152 chop-nod 1.480.02 – 0.330.05 1342223625 55745.669 152 chop-nod 1.450.02 – 0.270.05 1342223690 55746.726 152 chop-nod 1.510.02 – 0.310.04 1342223885 55752.614 152 chop-nod 1.510.03 – 0.260.05 1342223925 55753.611 152 chop-nod 1.580.03 – 0.260.06 R LMi 1342254139 56230.736 52 scan map 11.720.02 – 2.450.05 R UMa 1342221940 55712.986 52 scan map 2.810.05 – 0.470.05 1342246162 56073.792 52 scan map 2.990.04 – 0.560.04 1342246187 56074.789 52 scan map 3.120.06 – 0.590.05 1342246747 56085.991 52 scan map 3.270.06 – 0.690.04 1342246935 56088.951 52 scan map 3.330.06 – 0.570.02 RR Eri 1342190602 55239.380 152 chop-nod 1.750.02 – 0.350.05 RR UMi 1342254046 56229.690 52 scan map 2.970.01 – 0.530.02 RT Cnc 1342253343 56211.136 52 scan map 2.300.07 – 0.430.04 1342253469 56215.813 52 scan map 2.440.06 – 0.560.06 RU Her 1342190522 55239.065 152 chop-nod 8.370.03 – 1.430.03 1342191825 55264.384 152 chop-nod 8.220.03 – 1.380.04 1342228878 55822.857 152 chop-nod 6.610.03 – 1.140.03 ST UMa 1342247296 56101.709 52 scan map 1.630.03 – 0.320.09 1342247425 56103.683 52 scan map 1.630.04 – 0.190.06 SV Peg 1342220943 55696.010 152 chop-nod 13.030.05 – 1.870.16 1342221173 55697.791 152 chop-nod 13.170.04 – 1.830.12 1342222726 55729.050 152 chop-nod 12.880.05 – 1.840.11 • Notes.
+Appendix C UV excesses and dust attenuation
+[b] Source NUV FUV NUV FUV NUV FUV (Obs.) (Obs.) (SED) (SED) (excess) (excess) (Obs.) (Jy) (Jy) (Jy) (Jy) () AT Dra 2190 177 1700 7.6e 1.29 2.30 0.080 – – 0.080 BC CMi(∗) 1013 55 720 2.9 1.41 1.90 0.054 – – 0.054 BD Cam(∗) 5970 2330 4200 3.2 1.42 7.28 0.390 – – 0.390 BY Boo 3140 214 2100 8.3 1.50 2.58 0.068 – – 0.068 CG UMa 1498 108 1200 4.9 1.25 2.20 0.072 – – 0.072 DF Leo 590 36 1000 7.6 0.59 0.47 0.061 – – 0.061 EY Hya 110 82 23 2.8 4.8 290 0.740 0.99 1.67 1.461 FH Vir 400 49 370 1.4 1.08 3.5 0.120 0.14 0.24 0.133 IN Hya 620 44 1200 8.6 0.52 0.51 0.071 0.11 0.19 0.077 ome Vir(∗) 2950 179 2100 9.3 1.40 1.92 0.061 – – 0.061 R LMi(b) 40 2.4 7.6 5.2 5 46 0.040 4.50 7.63 0.9 R UMa 90 50 30 5.3 3.0 94 0.6 1.13 1.91 1.3 RR Eri 430 28 300 8.4 1.4 3.3 0.064 0.08 0.14 0.068 RR UMi 7200 1870 4100 1.8 1.76 1.04 0.26 – – 0.26 RT Cnc 190 32 170 3.7 1.12 9 0.19 0.76 1.29 0.32 RU Her(b) 80 5.75 27 5.8 3.0 10 0.08 0.56 0.95 0.12 RW Boo 410 23 180 5.3 2.3 4.3 0.057 0.17 0.29 0.064 RZ UMa(b) 18 8 22 3.6 0.82 22 0.4 0.51 0.86 0.6 ST UMa(∗) 994 168 680 2.7 1.46 6.2 0.17 0.25 0.43 0.20 SV Peg(b) 29 5.75 39 4.6 0.74 13 0.21 0.55 0.91 0.30 T Dra(a,b) 32 7.9 2.6 4.2 12.3 190 0.25 7.76 8.05 0.33 TU And(b) 15 5.75 19 4.3 0.8 13 0.96 0.98 1.67 1.91 UY Leo 150 27 29 7.6 5.2 36 0.18 0.56 0.95 0.27 V Eri 100 67 34 4.0 2.9 170 0.67 1.41 2.39 1.8 VY UMa 140 8 110 3.0 1.3 2.7 0.06 0.69 0.72 0.06 W Peg 180 13 41 5.8 4.4 22 0.065 0.98 1.67 0.13 Y CrB(b) 12 5.75 16 3.3 0.8 17 0.67 0.70 1.19 1.09 Y Gem 6000 5000 4.6 1.0 1300 5000 0.86 0.08 0.14 0.91 Z Cnc 73 12 9.2 2.4 7.9 5.0 0.17 0.28 0.48 0.21 • Notes.
 Column (1): Source, Col.
-(2): Obs ID, Col.
-(3): OBs date, Col.
-(4): exposure duration, Col.
-(5): observational mode, Col.
-(6): Flux in band B, Col.
-(7): Flux in band G, Col.
-(8): Flux in band R.
-Herschel/PACS B, G and R filters have effective wavelengths of 72, 103 and 167 m, and effective spectral widths of 21, 31, and 70 m respectively.
-[b] Source OBs ID OBs date duration mode Flux band B Flux band G Flux band R (MJD) (s) (Jy) (Jy) (Jy) T Dra 1342187357 55165.608 161 chop-nod 8.470.02 – 1.660.03 1342190576 55239.267 152 chop-nod 7.760.02 – 1.550.03 1342222753 55729.353 152 chop-nod 8.660.02 – 1.620.03 1342232432 55882.643 152 chop-nod 10.000.02 – 2.080.03 1342232465 55883.675 152 chop-nod 11.100.02 – 2.020.03 1342233065 55891.711 152 chop-nod 11.130.02 – 2.020.03 1342233205 55892.709 152 chop-nod 11.020.02 – 2.030.03 1342233356 55896.707 152 chop-nod 11.040.02 – 1.990.03 1342233587 55899.624 152 chop-nod 11.070.02 – 1.980.03 1342234216 55900.630 152 chop-nod 10.930.02 – 2.000.03 1342240701 55992.767 152 chop-nod 9.090.02 – 1.690.03 1342241298 55999.595 152 chop-nod 8.930.02 – 1740.03 V Eri 1342189943 55227.184 283 scan map 12.610.04 – 1.900.16 1342189944 55227.188 283 scan map 12.610.04 – 1.900.16 VY UMa 1342210606 55530.947 881 scan map 1.200.04 – 0.190.02 1342210607 55530.958 881 scan map 1.200.04 – 0.190.02 1342243219 56008.669 733 scan map – 0.640.02 0.230.03 1342243220 56008.678 733 scan map – 0.640.02 0.230.03 W Peg 1342222729 55729.060 152 chop-nod 6.610.03 – 1.280.04 Z Cnc 1342251911 56200.817 52 scan map 1.330.01 – 0.200.13
-Appendix D UV excesses and dust attenuation
+(2): Averaged observed NUV flux, Col.
+(3): Averaged observed FUV flux, Col.
+(4): NUV flux expected from SED modelling, Col.
+(5): FUV flux expected from SED modelling, Col.
+(6): NUV excess, Col.
+(7): FUV excess, Col.
+(8): Averaged ratio between FUV and NUV fluxes, Col.
+(9): Optical depth at NUV effective wavelength, Col.
+(10): Optical depth at FUV effective wavelength, Col.
+(11): Averaged ratio between FUV and NUV fluxes corrected from dust attenuation.
 We computed synthetic stellar photometry on the GALEX bands and compared them with the observed fluxes to estimate the UV excesses (similarly to Sahai et al.
@@
 Therefore, the stellar fluxes were estimated by extrapolating the input spectra following a blackbody distribution and convolving it with the transmission curves of their respective filters.
+Table 3 shows the observed and modelled properties of the UV emission for the sources used in this sample.
 We acknowledge that the real stellar spectrum differs from a blackbody approximation mainly due to the presence of (i) molecular absorption bands, and (ii) emission-line features due to possible chromospheric emission, that can produce a significant effect on the shape of the stellar spectra in the UV.
@@
 We highlight that these correlations can be seen despite the scatter produced by UV emission variability, and they become more pronounced when corrected for dust attenuation.
-[b] Source NUV FUV NUV FUV NUV FUV (Obs.) (Obs.) (SED) (SED) (excess) (excess) (Obs.) (Jy) (Jy) (Jy) (Jy) () AT Dra 2190 177 1700 7.6e 1.29 2.30 0.080 – – 0.080 BC CMi(∗) 1013 55 720 2.9 1.41 1.90 0.054 – – 0.054 BD Cam(∗) 5970 2330 4200 3.2 1.42 7.28 0.390 – – 0.390 BY Boo 3140 214 2100 8.3 1.50 2.58 0.068 – – 0.068 CG UMa 1498 108 1200 4.9 1.25 2.20 0.072 – – 0.072 DF Leo 590 36 1000 7.6 0.59 0.47 0.061 – – 0.061 EY Hya 110 82 23 2.8 4.8 290 0.740 0.99 1.67 1.461 FH Vir 400 49 370 1.4 1.08 3.5 0.120 0.14 0.24 0.133 IN Hya 620 44 1200 8.6 0.52 0.51 0.071 0.11 0.19 0.077 ome Vir(∗) 2950 179 2100 9.3 1.40 1.92 0.061 – – 0.061 R LMi(b) 40 2.4 7.6 5.2 5 46 0.040 4.50 7.63 0.9 R UMa 90 50 30 5.3 3.0 94 0.6 1.13 1.91 1.3 RR Eri 430 28 300 8.4 1.4 3.3 0.064 0.08 0.14 0.068 RR UMi 7200 1870 4100 1.8 1.76 1.04 0.26 – – 0.26 RT Cnc 190 32 170 3.7 1.12 9 0.19 0.76 1.29 0.32 RU Her(b) 80 5.75 27 5.8 3.0 10 0.08 0.56 0.95 0.12 RW Boo 410 23 180 5.3 2.3 4.3 0.057 0.17 0.29 0.064 RZ UMa(b) 18 8 22 3.6 0.82 22 0.4 0.51 0.86 0.6 ST UMa(∗) 994 168 680 2.7 1.46 6.2 0.17 0.25 0.43 0.20 SV Peg(b) 29 5.75 39 4.6 0.74 13 0.21 0.55 0.91 0.30 T Dra(a,b) 32 7.9 2.6 4.2 12.3 190 0.25 7.76 8.05 0.33 TU And(b) 15 5.75 19 4.3 0.8 13 0.96 0.98 1.67 1.91 UY Leo 150 27 29 7.6 5.2 36 0.18 0.56 0.95 0.27 V Eri 100 67 34 4.0 2.9 170 0.67 1.41 2.39 1.8 VY UMa 140 8 110 3.0 1.3 2.7 0.06 0.69 0.72 0.06 W Peg 180 13 41 5.8 4.4 22 0.065 0.98 1.67 0.13 Y CrB(b) 12 5.75 16 3.3 0.8 17 0.67 0.70 1.19 1.09 Y Gem 6000 5000 4.6 1.0 1300 5000 0.86 0.08 0.14 0.91 Z Cnc 73 12 9.2 2.4 7.9 5.0 0.17 0.28 0.48 0.21 • Notes.
-Column (1): Source, Col.
-(2): Averaged observed NUV flux, Col.
-(3): Averaged observed FUV flux, Col.
-(4): NUV flux expected from SED modelling, Col.
-(5): FUV flux expected from SED modelling, Col.
-(6): NUV excess, Col.
-(7): FUV excess, Col.
-(8): Averaged ratio between FUV and NUV fluxes, Col.
-(9): Optical depth at NUV effective wavelength, Col.
-(10): Optical depth at FUV effective wavelength, Col.
-(11): Averaged ratio between FUV and NUV fluxes corrected from dust attenuation.
+Appendix D SEDs and best-fit models for the rest of dusty AGB stars
+Fig. 14 shows the SEDs and their best-fit DUSTY models for the rest of dusty AGB stars to complement the sources already shown in Fig. 1.
 Appendix E Parameter uncertainties and correlations from the SED modelling
@@
 Therefore, it is likely the presence of some systematic differences between our estimates and NESS.
-On the other hand, anticorrelates with , this correlation, although it presents some scatter, which can be related with UV emission variability, and it is not corrected from dust attenuation.
+On the other hand, anticorrelates with , although it presents some scatter that can be related with UV emission variability and it is not corrected from dust attenuation.
 Finally, we crossmatched the NESS catalogue with GALEX and found that among the 485 AGB stars in NESS, 163 were detected in the NUV band and 95 were also detected in the FUV band.